# A multiwavelength ALMA view of gas and dust in binary protoplanetary system AS 205: Evidence of dust asymmetric distribution

### v1

@@ Abstract
We present Atacama Large Millimeter/Submillimeter Array observations of multi-wavelength dust emissions at 3.1 mm and 1.3 mm; along with molecular line emissions of CO(2–1), CO(3–2), 13CO(3–2), and C18O(3–2) at spatial resolutions of 7–45 AU towards the protoplanetary system AS 205. The dust emissions exhibit two distinct components of AS 205 N and AS 205 S, separated by 1.3 arcsec. While gas kinematics within the dust disk regions are dominated by Keplerian rotation, the more extended gas emission displays complex morphology and kinematics strongly affected by the binary gravitational interaction in the outer regions. The stellar masses of AS 205 N and AS 205 S are estimated at $0.78\pm 0.19$ M and $1.93\pm 0.86$ M, respectively. Azimuthal variation is observed in the spectral index distribution of both disks. In AS 205 N, the spectral index minimum in the southwest is coincident with the peaks of CO(212-1), CO(323-2), and 13CO(323-2) integrated intensity and the relative position of its southern counterpart. On the other hand, the spectral index distribution in AS 205 S exhibits two prominent maxima, with the one in the northeast aligning with the peak of 13CO(323-2), and the peak in the south coinciding with local maxima in CO(212-1) and CO(323-2) azimuthal profiles. These results suggest a correlation between dust grain size and/or optical depth with the gas distributions. Dust-trapping along the spiral arms possibly contribute to the spectral index minima in AS 205 N; however, the observed asymmetry across both disks suggests the involvement of additional mechanisms.
@@ I Introduction
The investigations conducted by Salyk et al. (2014) employed ALMA observations of $\rm{CO}(2-1)$  and its isotopologues at an angular resolution of 0.6′′, unveiled non-Keplerian extended emissions originating from the northern component. They are attributed to the effects of tidal interaction, disk wind, or a combination thereof. Similarly, Kurtovic et al. (2018) scrutinized the spiral-like features observed in the $\rm{CO}(2-1)$ line emanating from the northern component. Utilizing a higher angular resolution of 0.1′′, they proposed that these emissions could potentially arise from a fly-by scenario. Furthermore, Weber et al. (2023) later employed the same $\rm{CO}(2-1)$ line data to estimate the dynamic masses of AS 205 N and AS 205 S, determining them at 0.58 M and 0.42 M, respectively. Additionally, the authors concluded that the stars within the binary system are not gravitationally bound, thereby suggesting the plausibility of a fly-by scenario within the system.
@@ II Observations and data reduction
The Band 3 observations were conducted in six execution blocks (EBs). The correlator was configured in the Time-Division Multiplexing (TDM) mode, setting up both spectral windows exclusively to target broadband continuum emission. The Band 7 observations were carried out in two EBs. One was set to target the molecular lines of HCO+ (4–3), HCN (4–3) and $\rm{CO}(3-2)$; while the other covered $^{13}\rm{CO}(3-2)$, C18O (3–2), SO (8–7) and CN (3–2). Additionally, Band 7 continuum emission was captured in two spectral windows centered at 342.087 GHz and 343.495 GHz, each with a total bandwidth of 1.875 GHz. Detailed information regarding the observing parameters of these observations (i.e., observing time, antenna configuration, number of effective antennas, baseline, total time on source, and calibrators of each EB) is provided in Table 1. For both bands, the calibrated measurement sets were acquired upon request from the ALMA East Asia ARC and subsequently processed using CASA version 6.6.0.
@@ III.1 Dust emissions
The two well-separated disks in the 3.1 mm continuum image are fitted with a 2D Gaussian model using the CASA task imfit. For AS 205 N, we utilize only the SB dataset, which has a beam size of $275\times 188$ mas2, to mitigate any asymmetric features. Additionally, for the outer dust ring of AS 205 S, we fit an ellipse function to the maximum emission points using the routine MCMC based on the package emcee (Foreman-Mackey et al., 2013). Results of the best-fit values are presented in Table 3, where they are compared with the values derived from fitting the 1.3 mm data by Kurtovic et al. (2018).
@@ IV.1 Dust disk morphology and dust distribution
Figure 5: Top: Azimuthal distribution of the spectral index of AS 205 N within the radial distance of $0.2-0.4^{\prime\prime}$ (left) and the outer ring of AS 205 S, which is defined at $0.2-0.4^{\prime\prime}$ (right). Bottom: Azimuthal distribution of the normalized integrated intensity for CO(212-1), CO(323-2), and 13CO(323-2), within the radial distance of $0.8^{\prime\prime}$ in AS 205 N (left) and $0.6^{\prime\prime}$ in AS 205 S (right). The azimuthal angle $\psi$ is measured eastward from the north direction.
@@ IV.2 Possible origins of the azimuthal asymmetry in the spectral index distribution
IV.3 Gravitationally bound binary or Hyperbolic fly-by?
Gas emission in the AS 205 binary system shows strong connections and extended structures between the two components. While the observations do not reveal a circumbinary disk or ring, the gas emission exhibits a significant asymmetry, specifically with extended non-Keplerian gas at the velocity of approximately $5-7\,\rm km\,s^{-1}$ concentrated in the northern hemisphere. This non-Keplerian gas was first reported by Pontoppidan et al. (2011) and Brown et al. (2013) using spectrographic observations. Subsequent high-resolution mapping using ALMA by Salyk et al. (2014) and Kurtovic et al. (2018) confirmed this motion. The nature of this motion is currently under debate, with proposed mechanisms including a molecular disk wind or outflow originating from the disk, or a tidal interaction scenario, such as a hyperbolic fly-by.
A critical question related to this problem is whether the AS 205 system is gravitationally bound or undergoing a hyperbolic fly-by, which has been addressed previously by Weber et al. (2023). By comparing the disk components’ relative velocity with the system’s escape velocity, the authors proposed that the AS 205 system is undergoing a hyperbolic fly-by, independent of the line-of-sight distance of the system. However, their analysis is also strongly dependent on the constrained total stellar mass and plane-of-sky velocity of the system, which are substantially different from our result as estimated in Section III.2.2. Therefore, as an adaptation to their work, we re-perform the analysis as follows.
@@ IV.3 Gravitationally bound binary or Hyperbolic fly-by?
Figure 8: Adapted from Figure 11 in Weber et al. (2023), comparison of relative velocity ($v_{rel}$) between AS 205 N and AS 205 S to the system’s escape velocity $v_{esc}$, depending on its total mass and the components’ separation along the line-of-sight (los). The velocity difference is given as a multiple of the uncertainty of the relative velocity $\sigma=0.37\,\rm km\,s^{-1}$. The grey dashed line shows where the velocities are equal. Parameter space in the top-left is for a bound system, while the bottom-right is unbound, indicating a hyperbolic fly-by. The lower dashed line marks the total stellar mass estimate using 2D Keplerian fitting to the velocity map of CO(212-1) in Weber et al. (2023), the middle dashed line marks the result done with spectroscopic energy distribution (SED) fitting from Eisner et al. (2005) and Andrews et al. (2018), and the upper dashed line marks our result using PV diagram fitting to different CO isotopologues.
@@ V Summary
  • Our comparison analysis of the disk components’ relative velocity with the system’s escape velocity suggests that the mystery regarding the dynamical state of AS 205 remains an open question, unlike the concluded fly-by scenario as in Weber et al. (2023).
@@ Acknowledgment
We thank the anonymous reviewer for very constructive comments that help us to improve the quality of our manuscript; ALMA East Asia ARC, especially Dr. Atsushi Miyazaki, for their help in calibrating the data from the observations of projects 2015.1.00168.S and 2018.1.01198.S; and Dr. Akimasa Kataoka (NAOJ) and Dr. Kiyoaki Doi (MPIA) for their helpful discussions regarding the dust asymmetric distribution. This research was funded by Vingroup Innovation Foundation (VINIF) under project code VINIF.2023.DA.057. A part of this work was carried out while we were associate members of the Simons Astrophysics Group at ICISE (Quy Nhon, Viet Nam). We acknowledge the financial support from the Simons Foundation (No. 916424, N.H.) and the warm hospitality of the ICISE staff. Additionally, Nguyen Thi Phuong acknowledges financial support from the World Laboratory, Nguyen Tat Thang acknowledges support through a Neighbourhood, Development and International Cooperation Instrument (NDICI) scholarship funded by the European Union in the framework of the Erasmus+, Erasmus Mundus Joint Master in Astrophysics and Space Science – MASS. Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the granting authority, European Education and Culture Executive Agency (EACEA). Neither the European Union nor the granting authority can be held responsible for them. This paper makes use of the following ALMA data: ADS/JAO.ALMA#2015.1.00168.S, 2016.1.00484.L, and 2018.1.01198.S. ALMA is a partnership of ESO (representing its member states), NSF (USA), and NINS (Japan), together with NRC (Canada), NSTC and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO, and NAOJ.

### v2

@@ Abstract
We present Atacama Large Millimeter/Submillimeter Array observations of multi-wavelength dust emissions at 3.1 mm and 1.3 mm; along with molecular line emissions of CO(2–1), CO(3–2), 13CO(3–2), and C18O(3–2) at spatial resolutions of 7–45 AU towards the protoplanetary system AS 205. The dust emissions exhibit two distinct components of AS 205 N and AS 205 S, separated by 1.3 arcsec. While gas kinematics within the dust disk regions are dominated by Keplerian rotation, the more extended gas emission displays complex morphology and kinematics strongly affected by the binary gravitational interaction in the outer regions. The stellar masses of AS 205 N and AS 205 S are estimated at $0.78\pm 0.19$ and $1.93\pm 0.86$ M, respectively. Azimuthal variation is observed in the spectral index distribution of both disks. In AS 205 N, the spectral index minimum in the southwest is coincident with the peaks of CO(212-1), CO(323-2), and 13CO(323-2) integrated intensity and the relative position of its southern counterpart. On the other hand, the spectral index distribution in AS 205 S exhibits two prominent maxima, with the one in the northeast aligning with the peak of 13CO(323-2), and the peak in the south coinciding with local maxima in CO(212-1) and CO(323-2) azimuthal profiles. These results suggest a correlation between dust grain size and/or optical depth with the gas distributions. Dust trapping along the spiral arms possibly contributes to the spectral index minima in AS 205 N; however, the observed asymmetry across both disks suggests the involvement of additional mechanisms.
@@ I Introduction
The investigations conducted by Salyk et al. (2014) employed ALMA observations of $\rm{CO}(2-1)$  and its isotopologues at an angular resolution of 0.6′′, unveiled non-Keplerian extended emissions originating from the northern component. They are attributed to the effects of tidal interaction, disk wind, or a combination thereof. Similarly, Kurtovic et al. (2018) scrutinized the spiral-like features observed in the $\rm{CO}(2-1)$ line emanating from the northern component. Utilizing a higher angular resolution of 0.1′′, they proposed that these emissions could potentially arise from a flyby scenario. Furthermore, Weber et al. (2023) later employed the same $\rm{CO}(2-1)$ line data to estimate the dynamic masses of AS 205 N and AS 205 S, determining them at 0.58 M and 0.42 M, respectively. Additionally, the authors concluded that the stars within the binary system are not gravitationally bound, thereby suggesting the plausibility of a flyby scenario within the system.
@@ II Observations and data reduction
The Band 3 observations were conducted in six execution blocks (EBs). The correlator was configured in the Time-Division Multiplexing mode, setting up both spectral windows exclusively to target broadband continuum emission. The Band 7 observations were carried out in two EBs. One was set to target the molecular lines of HCO+ (4–3), HCN (4–3) and $\rm{CO}(3-2)$; while the other covered $^{13}\rm{CO}(3-2)$, C18O (3–2), SO (8–7) and CN (3–2). Additionally, Band 7 continuum emission was captured in two spectral windows centered at 342.087 GHz and 343.495 GHz, each with a total bandwidth of 1.875 GHz. Detailed information regarding the observing parameters of these observations (i.e., observing time, antenna configuration, number of effective antennas, baseline, total time on source, and calibrators of each EB) is provided in Table 1. For both bands, the calibrated measurement sets were acquired upon request from the East Asian ALMA Regional Center (EA-ARC) and subsequently processed using CASA version 6.6.0.
@@ III.1 Dust emissions
The two well-separated disks in the 3.1 mm continuum image are fitted with a 2D Gaussian model using the CASA task imfit. For AS 205 N, we utilize only the SB dataset, which has a beam size of $275\times 188$ mas2, to mitigate any asymmetric features. Additionally, for the outer dust ring of AS 205 S, we fit an ellipse function to the maximum emission points using the routine Markov-chained Monte-Carlo (MCMC) based on the package emcee (Foreman-Mackey et al., 2013). Results of the best-fit values are presented in Table 3, where they are compared with the values derived from fitting the 1.3 mm data by Kurtovic et al. (2018).
@@ IV.1 Dust disk morphology and dust distribution
Figure 5: Top: azimuthal distribution of dust continuum intensity at 3.1 and 1.3 mm. Middle: azimuthal distribution of the spectral index of AS 205 N within the radial distance of $0.2-0.4^{\prime\prime}$ (left) and the outer ring of AS 205 S, which is defined at $0.2-0.4^{\prime\prime}$ (right). Bottom: azimuthal distribution of the normalized integrated intensity for CO(212-1), CO(323-2), and 13CO(323-2), within the radial distance of $0.8^{\prime\prime}$ in AS 205 N (left) and $0.6^{\prime\prime}$ in AS 205 S (right). The azimuthal angle $\psi$ is measured eastward from the north direction.
@@ IV.2 Possible origins of the azimuthal asymmetry in the spectral index distribution
IV.3 Gravitationally bound binary or Hyperbolic flyby?
Gas emission in the AS 205 binary system shows strong connections and extended structures between the two components. While the observations do not reveal a circumbinary disk or ring, the gas emission exhibits a significant asymmetry, specifically with extended non-Keplerian gas at the velocity of approximately $5-7\,\rm km\,s^{-1}$ concentrated in the northern hemisphere. This non-Keplerian gas was first reported by Pontoppidan et al. (2011) and Brown et al. (2013) using spectrographic observations. Subsequent high-resolution mapping using ALMA by Salyk et al. (2014) and Kurtovic et al. (2018) confirmed this motion. The nature of this motion is currently under debate, with proposed mechanisms including a molecular disk wind or outflow originating from the disk, or a tidal interaction scenario, such as a hyperbolic flyby.
A critical question related to this problem is whether the AS 205 system is gravitationally bound or undergoing a hyperbolic flyby, which has been addressed previously by Weber et al. (2023). By comparing the disk components’ relative velocity with the system’s escape velocity, the authors proposed that the AS 205 system is undergoing a hyperbolic flyby, independent of the line-of-sight distance of the system. However, their analysis is also strongly dependent on the constrained total stellar mass and plane-of-sky velocity of the system, which are substantially different from our result as estimated in Section III.2.2. Therefore, as an adaptation to their work, we re-perform the analysis as follows.
@@ IV.3 Gravitationally bound binary or Hyperbolic flyby?
Figure 8: Adapted from Figure 11 in Weber et al. (2023), comparison of relative velocity ($v_{rel}$) between AS 205 N and AS 205 S to the system’s escape velocity $v_{esc}$, depending on its total mass and the components’ separation along the line-of-sight (los). The velocity difference is given as a multiple of the uncertainty of the relative velocity $\sigma=0.37\,\rm km\,s^{-1}$. The grey dashed line shows where the velocities are equal. Parameter space in the top-left is for a bound system, while the bottom-right is unbound, indicating a hyperbolic flyby. The lower dashed line marks the total stellar mass estimate using 2D Keplerian fitting to the velocity map of CO(212-1) in Weber et al. (2023), the middle dashed line marks the result done with spectroscopic energy distribution (SED) fitting from Eisner et al. (2005) and Andrews et al. (2018), and the upper dashed line marks our result using PV diagram fitting to different CO isotopologues.
@@ V Summary
  • Our comparison analysis of the disk components’ relative velocity with the system’s escape velocity suggests that the mystery regarding the dynamical state of AS 205 remains an open question, unlike the concluded flyby scenario as in Weber et al. (2023).
@@ Acknowledgment
We thank the anonymous reviewer for very constructive comments that helped us improve the quality of our manuscript; ALMA East Asian ARC, especially Dr. Atsushi Miyazaki, for their help in calibrating the data from the observations of projects 2015.1.00168.S and 2018.1.01198.S; and Dr. Akimasa Kataoka (NAOJ) and Dr. Kiyoaki Doi (MPIA) for their helpful discussions regarding the dust asymmetric distribution. This research was funded by Vingroup Innovation Foundation (VINIF) under project code VINIF.2023.DA.057. A part of this work was carried out while we were associate members of the Simons Astrophysics Group at ICISE (Quy Nhon, Viet Nam). We acknowledge the financial support from the Simons Foundation (No. 916424, N.H.) and the warm hospitality of the ICISE staff. Additionally, N.T.P. acknowledges financial support from the World Laboratory, N.T.T. acknowledges support through a Neighbourhood, Development and International Cooperation Instrument (NDICI) scholarship funded by the European Union in the framework of the Erasmus+, Erasmus Mundus Joint Master in Astrophysics and Space Science – MASS. Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the granting authority, European Education and Culture Executive Agency (EACEA). Neither the European Union nor the granting authority can be held responsible for them. This paper makes use of the following ALMA data: ADS/JAO.ALMA#2015.1.00168.S, 2016.1.00484.L, and 2018.1.01198.S. ALMA is a partnership of ESO (representing its member states), NSF (USA), and NINS (Japan), together with NRC (Canada), NSTC and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO, and NAOJ.